how many moles of helium would it take to fill a gas balloon with a volume of 1.0xEE6 liters at STP?
1
Expert's answer
2016-02-16T00:00:57-0500
pV=NRT
N=pv/RT=101.325*1e6/(8.314*273)=44 642 moles
Here we assume that temperature is 273 K (0 C).